Find an equatiin of the line having the given slope and containing the given point. M= -3,(5,9)
y=_____.

Answer by stanbon(75887)

Find an equatiin of the line having the given slope and containing the given point. M= -3,(5,9)
-----
Form: y = mx + b
Substitute and solve for "b":
9 = -3*5 + b
9 = -15 + b
b = 24
------
Equation:
y = -3x + 24
==================
Cheers,
Stan H.
==================